# Asymptotics for partitions over the Fibonacci numbers and related sequences

### 12 Dec 2023 v1

@@ Abstract.
In this paper, harkening back to ideas of Hardy and Ramanujan, Mahler and de Bruijn, with the addition of more recent results on the Fibonacci Dirichlet series, we determine the asymptotic number of ways $p_{F}(n)$ to write an integer as the sum of non-distinct Fibonacci numbers, as well as an asymptotic for the summatory function of the numbers $q_{F}(n)$—the number of ways to write $n$ as the sum of distinct Fibonacci numbers. This appears to be the first such asymptotic result concerning non-distinct partitions over Fibonacci numbers. As well, under weak conditions, we prove analogous results for a general linear recurrences.
@@ 1. Introduction
We let $p_{F}(n)$ denote the number of non-distinct partitions of $n$ over the Fibonacci sequence, that is, solutions to (1) in nonnegative integers $a_{k}$, and let $q_{F}(n)$ denote the number of distinct partitions of $n$ over the Fibonacci sequence, that is, solutions to (1) with $a_{k}\in\{0,1\}$. In recent work, Chow and Slattery [2] gave results on $q_{F}(n)$ (denoted $R(n)$ in their paper), and noted that their work shows that neither $q_{F}(n)$, nor its partial sums, have a ‘nice’ asymptotic formula. In particular, they showed that there is some oscillation in the partial sums of $q_{F}(n)$, and gave bounds on these oscillations. In this paper, we determine asymptotic formulas for both $p_{F}(n)$ and the partial sums of $q_{F}(n)$, and describe these oscillations.
@@
Let $p_{F}(n)$ and $q_{F}(n)$ be the number of partitions of $n$ over non-distinct and distinct Fibonacci numbers, respectively. Then, as $n\to\infty$,
$p_{F}(n)=\psi_{p_{F}}\left(\frac{\log n}{\log\varphi}\right)(\log n)^{-\frac{1}{2}}n^{\left(\frac{\log n}{2\log\varphi}+\frac{\log\log\varphi}{\log\varphi}+\frac{\log 5-\log\varphi+4\gamma}{2\log\varphi}-2\right)}\left(1+O\left(\frac{1}{\log n}\right)\right),$
and,
$\sum_{m\leqslant n}q_{F}(m)=\psi_{q_{F}}\left(\frac{\log n}{\log\varphi}\right)n^{\frac{\log 2}{\log\varphi}}\left(1+O\left(\frac{1}{n}\right)\right),$
where $\psi_{p_{F}}(x)$ and $\psi_{q_{F}}(x)$ are explicit positive $1$-periodic functions, $\varphi$ is the golden mean and $\gamma$ is the Euler–Mascheroni constant.
Theorem 1 seems to be the first asymptotic result chracterising non-distinct partitions over Fibonacci numbers. Distinct partitions over Fibonacci numbers have received considerable attention; see [2] and the references therein for details. In very recent work, addressing a question of Chow and Slattery [2], Kempton [5] has shown that $n^{-\frac{\log 2}{\log\varphi}}\sum_{m\leqslant n}q_{F}(m)$ is $\log$-periodic. Our proof of Theorem 1 gives an explicit description of this $\log$-periodic function.
This paper is organised as follows. In Sections 2 and 3, we focus on the case of $p_{F}(n)$ of Theorem 1. In particular, inspired by ideas of de Bruijn [1], we prove an exact formula for the generating function of $p_{F}(n)$ in Section 2. We then use a saddle-point method in Section 3 to give the asymptotic result for $p_{F}(n)$. In Section 4, we prove our main result for $q_{F}(n)$; this is done a bit more quickly and with less detail than the previous two sections since the methods and results are similar. Finally, in Section 5, we give the complete extension of Theorem 1 to the case of a linear recurrence with dominant root. In particular, suppose that $P_{k}$ is a strictly increasing linearly recurrent sequence of positive integers of degree at least $2$ with $P_{1}=1$, such that the characteristic polynomial $\chi_{P}(x)$ of $P_{k}$ has a single dominant root $\beta>0$ and $P_{k}=\lambda\beta^{k}+\lambda_{2}\beta_{2}^{k}+\cdots+\lambda_{r}\beta_{r}^{k},$ where $\lambda,\lambda_{2},\ldots,\lambda_{r}$ are constants and $\beta_{2},\ldots,\beta_{r}$ are the algebraic conjugates of $\beta$. In Section 5, we prove
@@
Let $p_{P}(n)$ and $q_{P}(n)$ be the number of partitions of $n$ over non-distinct and distinct members of the sequence $P_{k}$, respectively. Then, as $n\to\infty$,
$p_{P}(n)=\psi_{p_{P}}\left(\frac{\log n}{\log\beta}\right)(\log n)^{-\frac{1}{2}}n^{\left(\frac{\log n}{2\log\beta}+\frac{\log\log\beta}{\log\beta}+\frac{2\gamma}{\log\beta}-\frac{\log\lambda}{\log\beta}-\frac{3}{2}\right)}\left(1+O\left(\frac{1}{\log n}\right)\right),$
and for any fixed $\varepsilon\in(0,1)$,
$\sum_{m\leqslant n}q_{P}(m)=\psi_{q_{P}}\left(\frac{\log n}{\log\beta}\right)n^{\frac{\log 2}{\log\beta}}\left(1+O\left(n^{-\min\{{\log(\beta/|\beta_{2}|)}/{\log\beta},1-\varepsilon\}}\right)\right),$
where $\psi_{p_{P}}(x)$ and $\psi_{q_{P}}(x)$ are explicitly determinable positive $1$-periodic functions.
Remark 1.
Note that if $\beta>2$, then Theorem 2 implies the well-known result that the number of integers that have only ones and zeros in their base-$\beta$ expansion has density $\log 2/\log\beta<1$. This number is also the Hausdorff dimension of the subset of real numbers in the interval $[0,1]$ that have only ones and zeros in their base-$\beta$ expansions. $\Diamond$
@@ 2. An exact formula for for the generating series of $p_{F}(n)$
and $\zeta(1+z)=\zeta(1-4n)+\zeta^{\prime}(1-4n)(z+4n)+O((z+4n)^{2})$ along with the fact that $\zeta(1-4n)=B_{2n}/4n$; see Titchmarch [13, p. 19].
@@ 3. Fibonacci partitions $p_{F}(n)$ via the saddle point method
If $\Lambda(x)=\prod_{k\geqslant 1}\left(1\pm x^{\lambda_{k}}\right)^{\pm 1}$ generates a sequence $p_{\lambda}(n)$, and $s=\alpha$ is the solution of (13) with $F_{2}(e^{-s})$ replaced by $\Lambda(e^{-s})$, then, as $n$ tends to infinity,
@@
Here, the $\pm$ indicates the same choice in both instances, and $\alpha$ must be thought of as being replaced by its large-$n$ asymptotic expansion so that the asymptotic of $p_{\lambda}(n)$ represents a large-$n$ asymptotic.
@@
$n=\frac{-\log(\alpha\varphi e^{-c_{3}})}{\alpha\log\varphi}\left(1+O\left(\frac{\alpha}{\log\alpha}\right)\right).$
@@
(15) $=-\frac{\log\alpha}{\alpha\log\varphi}+\frac{\log(e^{c_{3}}/\varphi)}{\alpha\log\varphi}+O(1),$
@@
$\alpha=\frac{1}{n\log\varphi}\left(1+O\left(\frac{\log n}{n}\right)\right),$
@@
for all $x$ sufficiently large. Here, we have $x=\exp(n\log\varphi/\varphi e^{-c_{3}})$ and $y=1/\alpha\varphi e^{-c_{3}}$, so that
$\alpha\varphi e^{-c_{3}}=\frac{1}{W(\exp(n\log\varphi/\varphi e^{-c_{3}}))}\left(1+O\left(\frac{1}{n}\right)\right)=\frac{\varphi e^{-c_{3}}}{n\log\varphi}\left(1+O\left(\frac{\log n}{n}\right)\right),$
@@
$\log\alpha=-\log n-\log\log\varphi+O\left(\frac{\log n}{n}\right),$
and
$(\log\alpha)^{2}=(\log n)^{2}+2\log n\log\log\varphi+(\log\log\varphi)^{2}+O\left(\frac{(\log n)^{2}}{n}\right).$
@@
$\frac{2\pi j}{\log\varphi}\log\left(\tfrac{\alpha}{\sqrt{5}}\right)=-2\pi j\left(\frac{\log n}{{\log\varphi}}+\frac{\log\log\varphi+\log\sqrt{5}}{\log\varphi}\right)+O\left(\frac{\log n}{n}\right).$
@@
(20) $e^{n\alpha}=e^{n\cdot\frac{1}{n\log\varphi}\left(1+O\left(\frac{\log n}{n}\right)\right)}=e^{\frac{1}{\log\varphi}}\left(1+O\left(\frac{\log n}{n}\right)\right)$
@@
It remains to determine the small $\alpha$, and hence large $n$, asymptotics of $F_{2}(e^{-\alpha})$. To this end, we combine Theorem 3 with Proposition 7 to obtain, as $\alpha\to 0^{+}$,
@@
$=\frac{1}{2\log\varphi}\left((\log n)^{2}+2\log n\log\log\varphi+(\log\log\varphi)^{2}+O\left(\frac{(\log n)^{2}}{n}\right)\right)$
$\quad+\left(\log n+\log\log\varphi+O\left(\frac{\log n}{n}\right)\right)\left(\frac{c_{3}}{\log\varphi}-1\right)+c_{2}+2\gamma+\psi\left(\frac{\log n}{\log\varphi}\right)+O\left(\frac{1}{n}\right)$
$=\frac{(\log n)^{2}}{2\log\varphi}+(\log n)\left(\frac{\log\log\varphi}{\log\varphi}+\frac{c_{3}}{\log\varphi}-1\right)$
$\quad+\frac{(\log\log\varphi)^{2}}{2\log\varphi}+\left(\log\log\varphi\right)\left(\frac{c_{3}}{\log\varphi}-1\right)+c_{2}+2\gamma+\psi\left(\frac{\log n}{\log\varphi}\right)+O\left(\frac{(\log n)^{2}}{n}\right),$
@@
$F_{2}(e^{-\alpha})=\psi_{1}\left(\frac{\log n}{\log\varphi}\right)n^{\left(\frac{\log n}{2\log\varphi}+\frac{\log\log\varphi}{\log\varphi}+\frac{c_{3}}{\log\varphi}-1\right)}\left(1+O\left(\frac{(\log n)^{2}}{n}\right)\right),$
@@
$\psi_{1}\left(\frac{\log n}{\log\varphi}\right):=\exp\left(\frac{(\log\log\varphi)^{2}}{2\log\varphi}+\left(\log\log\varphi\right)\left(\frac{c_{3}}{\log\varphi}-1\right)+c_{2}+2\gamma+\psi\left(\frac{\log n}{\log\varphi}\right)\right).$
@@
4. The normalised mean of distinct Fibonacci partitions: the function $q_{F}(n)$
In this section, we deal completely with the case of distinct partitions over Fibonacci numbers, that is, the case of $q_{F}(n)$ in Theorem 1. The same method of the previous two sections is employed, so for convenience and brevity, assuming the reader is acquainted with those sections, we move a bit quicker here. Towards this goal, we let $q_{F}(n)$ be the number of ways to write the integer $n$ as the sum of distinct Fibonacci numbers (with only one 1). We define
$G_{2}(x):=\sum_{n\geqslant 0}\left(\sum_{m\leqslant n}q_{F}(m)\right)x^{n}=(1-x^{2})^{-1}\prod_{k\geqslant 1}\left(1+x^{F_{k}}\right)$
and consider the function
$G(e^{-s}):=\prod_{k\geqslant 1}\left(1+x^{F_{k}}\right).$
Here, we have that
[…]
where $a>0$, $\eta(z)=\sum_{m\geqslant 1}\frac{(-1)^{m+1}}{m^{z}}=(1-2^{1-z})\zeta(z)$ denotes the Dirichlet $\eta$-function. Note that $\eta(z)$ is entire and shares the trivial zeros of $\zeta(z)$. Considering values, we have that $\eta(1)=\log(2)$ and $\eta(0)=1/2$.
Theorem 5.
The function $G(x)$ defined above satisfies, as $s\to 0^{+}$,
$\log G(e^{-s})=(-\log s)\frac{\log 2}{\log\varphi}+\frac{\log 2(4\gamma+2\log 5-\log\varphi-\log 2)}{2\log\varphi}+\frac{s}{2}+f_{1}(s),$
where $f_{1}(s)$ satisfies $f_{1}(s)=f_{1}(\varphi s)+O(s^{2})$.
Proof.
As above, combining our knowledge of the functions $s^{-z},$ $\Gamma(z)$ and $\eta(1+z)$ with Proposition 1, the integrand $s^{-z}\Gamma(z)\eta(1+z)\zeta_{F}(z)$ has
  • a double pole at $z=0$,
  • a simple pole coming from $\Gamma(z)$ at $z=-1$,
  • double poles at $z\in-4\mathbb{N}$, and
where, as before the poles of $\Gamma(z)$ at $z=-(4m+2)$ for $m\in\mathbb{N}_{0}$ are cancelled by the corresponding zeros of $\zeta_{F}(z)$ and the poles of $\Gamma(z)$ at odd negative integers $z=-(2m+1)$ for $m\in\mathbb{N}$ are cancelled by the corresponding zeros of $\eta(1+z)$.
Using (5) and (6), Proposition 1, Lemma 1, and
$\eta(1+z)=\log 2+\left(\gamma\log 2-\frac{(\log 2)^{2}}{2}\right)z+O(z^{2}),$
we see that the double pole at $z=0$ contributes
$\underset{{{z}={0}}}{\textnormal{Res}}\left\{s^{-z}\Gamma(z)\eta(1+z)\zeta_{F}(z)\right\}=(-\log s)\frac{\log 2}{\log\varphi}+(4\gamma+2\log 5-\log\varphi-\log 2)\frac{\log 2}{2\log\varphi}.$
The simple pole at $z=-1$ contributes
$\underset{{{z}={-1}}}{\textnormal{Res}}\left\{s^{-z}\Gamma(z)\eta(1+z)\zeta_{F}(z)\right\}=s\,\eta(0)\,\zeta_{F}(-1)\cdot\underset{{{z}={-1}}}{\textnormal{Res}}\left\{\Gamma(z)\right\}=s\left(\frac{1}{2}\right)(-1)(-1)=\frac{s}{2}.$
Using the proof of Proposition 4, replacing the contribution $\zeta(1-4n)=B_{2n}/4n$ with the contribution $\eta(1-4n)=(1-2^{4n})\zeta(1-4n)=(1-2^{4n})B_{2n}/4n$, we have
$g_{1}(s):=\sum_{n=1}^{\infty}\underset{{{z}={-4n}}}{\textnormal{Res}}\left\{s^{-z}\Gamma(z)\eta(1+z)\zeta_{F}(z)\right\}=\sum_{n=1}^{\infty}\alpha_{n}\,s^{4n}-\log s\sum_{n=1}^{\infty}\beta_{n}s^{4n},$
where $\beta_{n}:={b_{-4n}\eta(1-4n)}/{(4n)!}$,
$\alpha_{n}:=\frac{(1-2^{4n})B_{2n}}{4n\cdot(4n)!}\left(b_{-4n}\left(\sum_{k=1}^{4n}\frac{1}{k}-\gamma\right)+c_{-4n}-\frac{5^{-2n}}{2}\frac{(4n)!}{((2n)!)^{2}}\right)+\frac{b_{-4n}}{(4n)!}\cdot\eta^{\prime}(1-4n),$
and $B_{2n}$ is the $2n$-th Bernoulli number. Moreover, as $s\to 0^{+}$, we have $g_{1}(s)=O(s^{2})$.
Finally, we have
$h_{1}(s):=$ $\,\sum_{k=0}^{\infty}\sum_{\begin{subarray}{c}n=-\infty\\ n\neq-k/2\end{subarray}}^{\infty}\underset{{{z}={s(n,k)}}}{\textnormal{Res}}\left\{s^{-z}\Gamma(z)\eta(1+z)\zeta_{F}(z)\right\}$
$=$ $\,\frac{1}{\log\varphi}\sum_{k=0}^{\infty}\sum_{\begin{subarray}{c}n=-\infty\\ n\neq-k/2\end{subarray}}^{\infty}(-1)^{k}{-s(n,k)\choose k}\left(\frac{s}{\sqrt{5}}\right)^{-s(n,k)}\Gamma(s(n,k))\,\eta(1+s(n,k)).$
Now, since $|(1-2^{it})|<2,$ for all real $t$, we have, from the result for $\zeta(1+it)$ that $\eta(1+it)=O(\log|t|)$, so that $\big{|}[k=0]h_{1}(s)\big{|}=O(1)$. For $k\neq 0$, we note that
$\big{|}\Gamma(s(n,k))\eta(1+s(n,k))\big{|}=\big{|}1-2^{-s(n,k)}\big{|}\cdot\big{|}\Gamma(s(n,k))\zeta(1+s(n,k))\big{|}<16^{k}\big{|}\Gamma(s(n,k))\zeta(1+s(n,k))\big{|}.$
Thus, in the proof of Proposition 5, we have the analogue of (12), with the added $16^{k}$ factor,
$\big{|}[k\neq 0]h_{1}(s)\big{|}<\frac{d}{\big{(}e^{\frac{\pi^{2}}{2\log\varphi}}-1\big{)}}\left(\frac{16s^{2}}{10\pi\cdot\log\varphi\cdot\big{(}e^{\frac{\pi^{2}}{2\log\varphi}}-1\big{)}}\right)^{k},$
which completes the proof of the theorem. The proof that $f_{1}(s):=g_{1}(s)+h_{1}(s)=f_{1}(\varphi s)+O(s^{2})$ follows mutatis mutandis the proof of Proposition 7. ∎
Corollary 2.
The function $G_{2}(x)$ defined above satisfies, as $s\to 0^{+}$,
$\log G_{2}(e^{-s})=(-\log s)\left(\frac{\log 2}{\log\varphi}+1\right)+\frac{\log 2(4\gamma+2\log 5-\log\varphi-\log 2)}{2\log\varphi}+2\gamma-\log 2+f_{1}(s)+O(s),$
where $f_{1}(s)$ satisfies $f_{1}(s)=f_{1}(\varphi s)+O(s^{2})$.
Proof.
We need only determine the asymptotics of $-\log(1-e^{-2s})$ and add those to the result in Theorem 5. To this end, note that $1-e^{-2s}=(1+e^{-s})(1-e^{-s})$. Now, as $s\to 0^{+}$, we have that $-\log(1+e^{-s})=-\log 2+O(s).$ To find the asymptotics of $-\log(1-e^{-s})$, we appeal to Lemma 2, which gives as $s\to 0^{+}$,
$-\log(1-e^{-s})=-\log s+2\gamma+O(s),$
so that $-\log(1-e^{-2s})=-\log s+2\gamma-\log 2+O(s),$ which when added to the statement in Theorem 5 yields the desired result. ∎
We find the saddle point as in the previous section. Here, we note that the saddle point $s=\alpha$ is the solution to
$\frac{d}{ds}\left(s+\frac{1}{n}\log G_{2}(e^{-s})\right)=0.$
Here, we use the asymptotics in Corollary 2 to give that
$n=-\left.\frac{d}{ds}\log G_{2}(e^{-s})\right|_{s=\alpha}=\frac{1}{\alpha}\left(\frac{\log 2}{\log\varphi}+1\right)+O(1)=\frac{1}{\alpha}\left(\frac{\log 2}{\log\varphi}+1\right)\left(1+O(\alpha)\right).$
Thus,
$\left.\frac{dn}{ds}\right|_{s=\alpha}=-\frac{1}{\alpha^{2}}\left(\frac{\log 2}{\log\varphi}+1\right)\left(1+O(\alpha^{2})\right).$
This gives the following result.
Proposition 8.
$\alpha=\frac{1}{n}\left(\frac{\log 2}{\log\varphi}+1\right)\left(1+O\left(\frac{1}{n}\right)\right),$
and
$\left.\frac{dn}{ds}\right|_{s=\alpha}=-\frac{n^{2}}{\left(\frac{\log 2}{\log\varphi}+1\right)}\left(1+O\left(\frac{1}{n^{2}}\right)\right).$
Corollary 3.
For sufficiently small $\alpha>0$, or equivalently, for sufficiently large $n$,
$-\log\alpha=\log n+\log\left(\frac{\log 2}{\log\varphi}+1\right)+O\left(\frac{1}{n}\right).$
We are now prepared to finish the proof of Theorem 1. Again, we apply Theorem 4.
Proof of Theorem 1 for $q_{F}(n)$.
As before, we evaluate the necessary pieces Theorem 4, with the appropriate generating function and saddle points. Proposition 8 gives, for any fixed $\varepsilon>0$,
(22) $e^{n\alpha}=e^{\left(\frac{\log 2}{\log\varphi}+1\right)}\left(1+O\left(\frac{1}{n}\right)\right)$
and
(23) $\left(\sqrt{\frac{1}{-\left.\frac{dn}{da}\right|_{a=\alpha}}}+O\left(\frac{1}{n^{3/2}}\right)\right)=\frac{1}{n}\sqrt{\frac{\log 2}{\log\varphi}+1}\left(1+O\left(\frac{1}{n^{2}}\right)\right).$
It remains to determine the small $\alpha$, and hence large $n$, asymptotics of $G_{2}(e^{-\alpha})$. To this end, we combine Corollary 2 with Corollary 3 to obtain, as $\alpha\to 0^{+}$, or, equivalently, as $n\to\infty$,
$\log G_{2}(e^{-\alpha})$ $=(\log n)\left(\frac{\log 2}{\log\varphi}+1\right)+\left(\frac{\log 2}{\log\varphi}+1\right)\log\left(\frac{\log 2}{\log\varphi}+1\right)$
$\qquad\qquad+\frac{\log 2(4\gamma+2\log 5-\log\varphi-\log 2)}{2\log\varphi}+2\gamma-\log 2+\psi_{2}\left(\frac{\log n}{\log\varphi}\right)+O\left(\frac{1}{n}\right),$
where $\psi_{2}(x)$ is an explicit $1$-periodic function. Thus, we have
$G_{2}(e^{-\alpha})=\psi_{3}\left(\frac{\log n}{\log\varphi}\right)n^{\left(\frac{\log 2}{\log\varphi}+1\right)}\left(1+O\left(\frac{1}{n}\right)\right),$
where $\psi_{3}(x)$ is an explicit $1$-periodic function. Combining this with equations (22) and (23) finishes the proof of the theorem. ∎

### 15 Dec 2023 v2

@@ Abstract.
In this paper, harkening back to ideas of Hardy and Ramanujan, Mahler and de Bruijn, with the addition of more recent results on the Fibonacci Dirichlet series, we determine the asymptotic number of ways $p_{F}(n)$ to write an integer as the sum of non-distinct Fibonacci numbers. This appears to be the first such asymptotic result concerning non-distinct partitions over Fibonacci numbers. As well, under weak conditions, we prove analogous results for a general linear recurrences.
@@ 1. Introduction
We let $p_{F}(n)$ denote the number of non-distinct partitions of $n$ over the Fibonacci sequence, that is, solutions to (1) in nonnegative integers $a_{k}$. In recent work, Chow and Slattery [2] gave results on the number distinct partitions, $q_{F}(n)$ over the Fibonacci sequence, and noted that their work shows that neither $q_{F}(n)$, nor its partial sums, have a ‘nice’ asymptotic formula. In particular, they showed that there is some oscillation in the partial sums of $q_{F}(n)$, and gave bounds on these oscillations. In a similar vein, very recently, Sana [12] showed that there are also oscillations in the partial sums of the powers $(q_{F}(n))^{m}$ for each $m$. In this paper, with a related motivation, we determine asymptotic formulas for $p_{F}(n)$ and describe the oscillations that occur.
@@
Let $p_{F}(n)$ be the number of partitions of $n$ over non-distinct Fibonacci numbers. Then, as $n\to\infty$,
$p_{F}(n)=\psi_{p_{F}}\left(\frac{\log n}{\log\varphi}\right)(\log n)^{-\frac{1}{2}}n^{\left(\frac{\log n}{2\log\varphi}+\frac{\log 5-\log\varphi+4\gamma}{2\log\varphi}-2+O\left(\frac{1}{\log n}\right)\right)}\left(1+O\left(\frac{1}{\log n}\right)\right),$
where $\psi_{p_{F}}(x)$ is an explicit positive $1$-periodic function, $\varphi$ is the golden mean and $\gamma$ is the Euler–Mascheroni constant.
Theorem 1 seems to be the first asymptotic result chracterising non-distinct partitions over Fibonacci numbers. Distinct partitions over Fibonacci numbers have received considerable attention; see [2] and the references therein for details. In very recent work, addressing a question of Chow and Slattery [2], Kempton [5] has shown that $n^{-\frac{\log 2}{\log\varphi}}\sum_{m\leqslant n}q_{F}(m)$ is $\log$-periodic. Our proof of Theorem 1 gives a description of the related $\log$-periodic function for $p_{F}(n)$.
This paper is organised as follows. In Sections 2 and 3, we focus the proof of Theorem 1. In particular, inspired by ideas of de Bruijn [1], we prove an exact formula for the generating function of $p_{F}(n)$ in Section 2. We then use a saddle-point method in Section 3 to give the asymptotic result for $p_{F}(n)$. Finally, in Section 4, we give the complete extension of Theorem 1 to the case of a linear recurrence with dominant root. In particular, suppose that $P_{k}$ is a strictly increasing linearly recurrent sequence of positive integers of degree at least $2$ with $P_{1}=1$, such that the characteristic polynomial $\chi_{P}(x)$ of $P_{k}$ has a single dominant root $\beta>0$ and $P_{k}=\lambda\beta^{k}+\lambda_{2}\beta_{2}^{k}+\cdots+\lambda_{r}\beta_{r}^{k},$ where $\lambda,\lambda_{2},\ldots,\lambda_{r}$ are constants and $\beta_{2},\ldots,\beta_{r}$ are the algebraic conjugates of $\beta$. In Section 4, we prove
@@
Let $p_{P}(n)$ be the number of partitions of $n$ over non-distinct elements of the sequence $P_{k}$. Then, as $n\to\infty$,
$p_{P}(n)=\psi_{p_{P}}\left(\frac{\log n}{\log\beta}\right)(\log n)^{-\frac{1}{2}}n^{\left(\frac{\log n}{2\log\beta}+\frac{2\gamma}{\log\beta}-\frac{\log\lambda}{\log\beta}-\frac{3}{2}+O\left(\frac{1}{\log n}\right)\right)}\left(1+O\left(\frac{1}{\log n}\right)\right),$
where $\psi_{p_{P}}(x)$ is an explicitly determinable positive $1$-periodic function.
@@ 2. An exact formula for for the generating series of $p_{F}(n)$
and $\zeta(1+z)=\zeta(1-4n)+\zeta^{\prime}(1-4n)(z+4n)+O((z+4n)^{2})$ along with the fact that $\zeta(1-4n)=B_{2n}/4n$; see Titchmarch [14, p. 19].
@@ 3. Fibonacci partitions $p_{F}(n)$ via the saddle point method
If $\Lambda(x)=\prod_{k\geqslant 1}\left(1-x^{\lambda_{k}}\right)^{-1}$ generates a sequence $p_{\lambda}(n)$, and $s=\alpha$ is the solution of (13) with $F_{2}(e^{-s})$ replaced by $\Lambda(e^{-s})$, then, as $n$ tends to infinity,
@@
Here, $\alpha$ must be thought of as being replaced by its large-$n$ asymptotic expansion so that the asymptotic of $p_{\lambda}(n)$ represents a large-$n$ asymptotic.
@@
$n=\frac{-\log(\alpha)}{\alpha\log\varphi}\left(1+O\left(\frac{1}{\log\alpha}\right)\right).$
@@
(15) $=-\frac{\log\alpha}{\alpha\log\varphi}+O\left(\frac{1}{\alpha}\right),$
@@
$\alpha=\frac{1}{n\log\varphi}\left(1+O\left(\frac{1}{\log n}\right)\right),$
@@
for all $x$ sufficiently large. Here, we have $x=\exp(n\log\varphi)$ and $y=1/\alpha$, so that
$\alpha=\frac{1}{W(\exp(n\log\varphi))}\left(1+O\left(\frac{1}{\log n}\right)\right)=\frac{1}{n\log\varphi}\left(1+O\left(\frac{1}{\log n}\right)\right),$
@@
$\log\alpha=-\log n-\log\log\varphi+O\left(\frac{1}{\log n}\right).$
@@
$\frac{2\pi j}{\log\varphi}\log\left(\tfrac{\alpha}{\sqrt{5}}\right)=-2\pi j\left(\frac{\log n}{{\log\varphi}}+\frac{\log\log\varphi+\log\sqrt{5}}{\log\varphi}\right)+O\left(\frac{1}{\log n}\right).$
@@
(20) $e^{n\alpha}=e^{n\cdot\frac{1}{n\log\varphi}\left(1+O\left(\frac{1}{\log n}\right)\right)}=e^{\frac{1}{\log\varphi}}\left(1+O\left(\frac{1}{\log n}\right)\right)$
@@
It remains to determine the small $\alpha$, and hence large $n$, asymptotics of $F_{2}(e^{-\alpha})$. To this end, we combine Theorem 3 with Corollary 1 and Proposition 7 to obtain, as $\alpha\to 0^{+}$,
@@
$=\frac{1}{2\log\varphi}(\log n)^{2}\left(1+O\left(\frac{1}{(\log n)^{2}}\right)\right)$
$\quad+\left(\log n+\log\log\varphi+O\left(\frac{1}{\log n}\right)\right)\left(\frac{c_{3}}{\log\varphi}-1\right)+c_{2}+2\gamma+\psi\left(\frac{\log n}{\log\varphi}\right)+O\left(\frac{1}{n}\right)$
$=\frac{(\log n)^{2}}{2\log\varphi}\left(1+O\left(\frac{1}{(\log n)^{2}}\right)\right)+(\log n)\left(\frac{c_{3}}{\log\varphi}-1\right)$
$\quad+\left(\log\log\varphi\right)\left(\frac{c_{3}}{\log\varphi}-1\right)+c_{2}+2\gamma+\psi\left(\frac{\log n}{\log\varphi}\right)+O\left(\frac{1}{(\log n)^{2}}\right),$
@@
$F_{2}(e^{-\alpha})=\psi_{1}\left(\frac{\log n}{\log\varphi}\right)n^{\left(\frac{\log n}{2\log\varphi}\left(1+O\left(\frac{1}{(\log n)^{2}}\right)\right)+\frac{c_{3}}{\log\varphi}-1\right)}\left(1+O\left(\frac{(\log n)^{2}}{n}\right)\right),$
@@
$\psi_{1}\left(\frac{\log n}{\log\varphi}\right):=\exp\left(\left(\log\log\varphi\right)\left(\frac{c_{3}}{\log\varphi}-1\right)+c_{2}+2\gamma+\psi\left(\frac{\log n}{\log\varphi}\right)\right).$